C2H4 + HCL --> C2H5CL dH=-72.3 kj find dE if 98g c2h4 and 109.5 g hcl is reacted at 300 K.

Grade:12

1 Answers

noel thekkekara
12 years ago

no  of moles of c2h2=?

molecular wt of c2h4=28

oso no of moles=98/28=3.5

so1mole of c2h4 reacts with 1 mole of HCl;3.5 of hcl will react

dH=3.5*72.3kj=253.05

dH=dU+d(nrt)

so calculate and get the answer